OpenStudy (anonymous):

Using the formula A=lw, find the are of the rectangle 8x^3 and 6x^2 I got 48x^6 Is that correct

OpenStudy (cwrw238):

48 is right but x^3 * x^2 = x^5 so its 48x^5

OpenStudy (anonymous):

\[\huge x^a \times x^b = x^{a+b}\]

OpenStudy (cwrw238):

you add the powers when multiplying

OpenStudy (anonymous):

So: \[\huge x^3 \times x^2 = x^{??}\]

OpenStudy (anonymous):

Thank you, I forgot to add instead of multiplying.